# London WC1E 6BT, United Kingdom

### v1

@@ 2.3 ALP decays
The interaction between the ALPs and active neutrinos in eq. (7) causes the ALP to decay. This decay channel occurs at tree level, thus dominating the ALP decay width, and it is given by
$\Gamma^{a\rightarrow\nu\nu}=\frac{1}{f_{a}^{2}}\frac{m_{N}^{2}m_{a}U_{eN}^{4}}{2\pi}\sqrt{1-\frac{4m_{\nu}^{2}}{m_{a}^{2}}}\left(1-\frac{2m_{\nu}^{2}}{m_{a}^{2}}\right)\simeq\frac{m_{N}^{2}m_{a}U_{eN}^{4}}{2\pi f_{a}^{2}}$ (25)
@@
No other decay channels are available at tree level. However, as it will be seen below (see section 4), most of the constraints on axion-like particles come from their interactions with electrons and photons. Though the Lagrangian in eq. (9) does not produce axion-electron or axion-photon interactions at tree level, such couplings can be induced at 1-loop and 2-loop respectively, as shown in Figure 2.
@@
$g_{a\gamma}$ $\approx\frac{e^{2}g_{ae}}{2\pi^{2}m_{e}}\left(1+\frac{1}{12}\frac{m_{a}^{2}}{m_{e}^{2}}\right)=\frac{\sqrt{2}e^{2}G_{F}|U_{eN}|^{4}m_{N}^{2}}{32\pi^{4}f_{a}}\left(1+\frac{1}{12}\frac{m_{a}^{2}}{m_{e}^{2}}\right),$ (26)
@@
Nevertheless, for light ALPs, $m_{a}\lesssim 1$ keV, which is the focus of this study, the only open decay channels for the ALPs are either a pair of active neutrinos or a pair of photons. But, as seen in Fig. 2, the diphoton decay happens at 2-loop and thus it is significantly suppressed with respect to the decay to active neutrinos. For example, the decay width to photons of a 1 keV ALP with decay constant of 1 TeV will be of order $10^{-46}$ GeV, negligible compared to decay width to active neutrinos in eq. (25).
@@
$\tau_{a}=1~{}\rm{sec}\times\left(\frac{1~{}\rm{GeV}}{m_{N}}\right)^{2}\times\left(\frac{1~{}\rm{keV}}{m_{a}}\right)\times\left(\frac{2.03\times 10^{-6}}{|U_{eN}|^{2}}\right)^{2}\times\left(\frac{f_{a}}{1~{}\rm{TeV}}\right)^{2}.$ (27)
@@ 3 Cosmological history of HNLs and ALPs
The change on the energy and number densities of the various particle species due to the expansion of the universe and the processes in Fig. 3 can be computed with the Boltzmann equations Gondolo:1990dk ; Escudero:2018mvt
@@
$\frac{dn_{X}}{dt}+3Hn_{X}$ $=\frac{\delta n_{X}}{\delta t}=\int g_{X}\frac{d^{3}p}{(2\pi)^{3}}\mathcal{C}[f],$ (28)
@@
$\gamma_{X\rightarrow Y}=\langle\sigma_{X\rightarrow Y}v\rangle n_{X}^{eq,2}=\langle\Gamma_{X\rightarrow Y}\rangle n_{X}^{eq},$ (29)
@@
$\gamma_{X\to Y_{1}\cdots}=n_{X}^{eq}(z)\frac{K_{1}(z)}{K_{2}(z)}\Gamma_{X},$ (30)
@@ 3.1 Abundances before BBN
We start by assuming that after some inflationary epoch and subsequent reheating, the SM particles are in equilibrium in the early Universe. The actual value of the reheating temperature is no importance, since most particles will be at their equilibrium densities, but it must be low enough to ensure that the ALPs are not strongly coupled after reheating. The HNLs are thus assumed to be in thermal and kinematical equilibrium with the SM, but the ALPs are not. The evolution of the densities of HNL and ALP can be described by coupled Boltzmann equations, as in eq. (28). Before neutrino decoupling, the HNL number density is given by
$\frac{dn_{N}}{dt}+3Hn_{N}=$ $-\left\langle\sigma_{NN\rightarrow SM}~{}v\right\rangle\left(n_{N}^{2}-n_{N}^{\mathrm{eq},2}\right)-\left\langle\sigma_{NN\to aa}~{}v\right\rangle\left(n_{N}^{2}-n_{N}^{\mathrm{eq},2}\frac{n_{a}^{2}}{n_{a}^{\mathrm{eq},2}}\right)$ (32)
@@
$\frac{dn_{a}}{dt}+3Hn_{a}$ $=-\left\langle\sigma_{aa\rightarrow\nu\nu}~{}v\right\rangle\left(n_{a}^{2}-n_{a}^{\mathrm{eq},2}\right)-\left\langle\sigma_{aa\to NN}~{}v\right\rangle\left(n_{a}^{2}-n_{a}^{\mathrm{eq},2}\frac{n_{N}^{2}}{n_{N}^{\mathrm{eq},2}}\right)$ (33)
@@
$-\gamma_{N\rightarrow SM}\left(\frac{Y_{N}}{Y_{N}^{\mathrm{eq}}}-1\right)-\gamma_{N\rightarrow a\nu}\left(\frac{Y_{N}}{Y_{N}^{\mathrm{eq}}}-\frac{Y_{a}}{Y_{a}^{\mathrm{eq}}}\right),$ (35)
@@
$-\gamma_{a\rightarrow\nu\nu}\left(\frac{Y_{a}}{Y_{a}^{\mathrm{eq}}}-1\right)+\gamma_{N\rightarrow a\nu}\left(\frac{Y_{N}}{Y_{N}^{\mathrm{eq}}}-\frac{Y_{a}}{Y_{a}^{\mathrm{eq}}}\right),$ (36)
@@ 3.2 Abundance of ALPs after BBN and temperature evolution
As we saw before, in all scenarios we consider, the abundance of HNLs depletes very fast before BBN. As a consequence, the main production mechanisms for ALPs, which were HNL-ALP scattering and HNL decays, are really inefficient after BBN and thus the ALP abundance only decreases at late times. Even the processes $aa\leftrightarrow N\nu$ and $aa\leftrightarrow\nu\nu$ fall out of thermal equilibrium significantly before BBN, as seen below in Fig. 4. Therefore, the abundance of ALPs after the end of BBN is solely determined by their decays. Their Boltzmann equation is therefore given byBoyarsky:2021yoh
$zHs\frac{dY_{a}}{dz}=-\gamma_{a\rightarrow\nu\nu}\left(\frac{Y_{a}}{Y_{a}^{\mathrm{eq}}}-1\right)\quad\quad\rm{with}~{}z>z_{\rm BBN}.$ (37)
@@
Before neutrino decoupling, the temperatures of the photon and neutrino baths evolved together, changing only via adiabatic cooling since ALP decays are negligible before BBN Hufnagel:2018bjp . After the end of BBN, however, the decays of ALPs to neutrinos, $a\to\nu\nu$ become relevant, which will introduce energy into the neutrino bath and thus increase its temperature. An increase on the neutrino temperature relative with the photon temperature at the time of recombination has a strong impact on the number of effective neutrino degrees of freedom $N_{\rm eff}$. Because of this, we need to track separately the evolution of the photon and neutrino temperatures from the end of BBN until the formation of the CMB. Neglecting chemical potentials and effects from neutrino oscillations, the evolution of the temperatures with time is given by Escudero:2018mvt ; EscuderoAbenza:2020cmq
@@
$\frac{dT_{\nu}}{dt}$ $=-\frac{12H\rho_{\nu}+3H(\rho_{a}+p_{a})+\delta\rho_{a}/\delta t}{3\tfrac{\partial\rho_{\nu}}{\partial T_{\nu}}+\tfrac{\partial\rho_{a}}{\partial T_{\nu}}}$ (38)
where $\rho_{i}$ and $p_{i}$ are the energy density and pressure of particle species $i$, and $\delta\rho_{a}/\delta t$ is the energy exchange rate from ALPs to neutrinos, given by the collision operator in eq.(28). Equation (38) takes into account the decoupling of the electrons from the photon bath when they become non-relativistic, as well as the decoupling of the ALP from the neutrino bath when $T_{\nu}\gtrsim m_{a}$.
@@ 4 Constraints from cosmology, astrophysics and direct searches
In this work we have introduced two new particles, a HNL and an ALP that modify very significantly the evolution of the early Universe compared to $\Lambda$CDM. The cosmological and astrophysical implications of both of these species are well known separately Cadamuro:2011fd ; Marsh:2015xka ; Drewes:2015iva . However, their combination has a somewhat different effect, as the HNLs are allowed to be lighter than in most previous scenarios, due to the weakening of the BBN bound (see below), and also due to the ALPs coupling exclusively to neutrinos via the HNL portal. Consequently we need to evaluate the predictions of our model towards known cosmological and astrophysical observables and hence assess whether their strong constraints render the model invalid or there are still parameter combinations that are unconstrained. For this purpose we study how this model affects the formation of primordial elements (BBN), the observations of the CMB, as well as astrophysical constraints such as those from the observation of the supernova SN1987A, and others such as extra-galactic background light (EBL) or X-ray constraints. Additionally, the new decay channel for HNLs available in this model modifies the prospects of direct searches for HNLs at colliders, fixed target or beam dump experiments. Consequently, we also study the sensitivity of our model to direct searches for HNLs at future facilities.
@@ 4.1 Big Bang Nucleosynthesis
The predictions of $\Lambda$CDM with regards to the formation of the primordial elements match very well with their observed present abundances [^2] The present-day abundance of ${}^{7}Li$ is currently in disagreement with the predictions from standard cosmology. Attempts have been made to provide an explanation in modified cosmologies, with varying degrees of success Depta:2020zbh ; Balazs:2022tjl . Therefore, it remains unclear whether this is caused by modifications over $\Lambda$CDM or an inaccurate measurement of their present abundance. Iocco:2008va ; Coc:2017pxv . These predictions are, however very sensitive to the cosmological state of the Universe at any time between neutrino decoupling and the CMB. Any modification of the temperature of neutrino decoupling, the rate of expansion of the Universe or energy injection in the primordial plasma, may cause a catastrophic change on the formation of light elements. This is specially true in our scenario, where the decays of HNLs into mesons disturb the $p\leftrightarrow n$ conversion processes that set the initial proton and neutron abundances for BBN Boyarsky:2020dzc . If the HNLs decay fast enough, however, the abundances of protons and neutrons have time to restore to the expected values from $\Lambda$CDM, and thus there is no effect on BBN. Hence, we set a conservative limit on the HNL lifetime of $\tau_{N}<0.023$Boyarsky:2020dzc , beyond which the decays will disturb the BBN abundances too much and is therefore marked as excluded in our scenario.
@@
In addition to the decays of the HNLs, the decays of ALPs may also modify the primordial plasma and affect BBN. If the ALPs decay before neutrino decoupling, since their primary decay channel is $a\to\nu\nu$, they would modify the neutrino spectrum and increase the temperature of the neutrino bath, thereby delaying neutrino decoupling and the formation of primordial elements. Consequently, we do not consider scenarios where the ALPs have short lifetimes, and to ensure that they do not directly affect BBN, we only consider ALPs decaying after the end of BBN, i.e $\tau_{a}>10^{4}$ s. Subsequent decays of the ALPs may also dissociate the formed elements and modify their abundance Hufnagel:2018bjp , however this dissociation is mostly caused by electromagnetic cascades, which are rare in our scenario since the decay rate of $a\to\gamma\gamma$ is negligible.
@@
Even if the ALPs do not decay before BBN, they may still affect the formation of the primordial elements if they are too abundant. Since ALPs are relativistic before neutrino decoupling, a sufficient abundance would act as dark radiation and thus modify the Hubble rate during radiation domination. This effect can be understood as increase in the neutrino temperature and therefore results in a modification on the value of $N_{\rm eff}$ at the time of BBN GAMBITCosmologyWorkgroup:2020htv . Since the contribution from dark radiation is only relevant when the ALPs are close to thermal equilibrium, we can parametrize the deviation from $\Lambda$CDM as
$\Delta N_{\rm eff}^{\rm BBN}\approx\frac{\rho_{a}}{\rho_{\gamma}}\approx\frac{\rho_{a}^{eq}n_{a}}{\rho_{\gamma}n_{a}^{eq}}$ (39)
@@
where $\rho_{\gamma}$ is the energy density in photons at BBN. Though $N_{\rm eff}$ is not directly measured at BBN, it can inferred from the measured value at recombination by Planck Planck:2018vyg , to be Fields:2019pfx $N_{\rm eff}^{\rm BBN}=2.86\pm 0.15$, hence we set an upper bound on the ALP abundance at BBN by requiring $\Delta N_{\rm eff}^{\rm BBN}\lesssim 0.2$.

### v2

@@ 2.3 ALP decays
The interaction between the ALPs and active neutrinos in eq. (2.7) causes the ALP to decay. This decay channel occurs at tree level, thus dominating the ALP decay width, and it is given by
$\Gamma^{a\rightarrow\nu\nu}=\frac{1}{f_{a}^{2}}\frac{m_{N}^{2}m_{a}U_{eN}^{4}}{2\pi}\sqrt{1-\frac{4m_{\nu}^{2}}{m_{a}^{2}}}\left(1-\frac{2m_{\nu}^{2}}{m_{a}^{2}}\right)\simeq\frac{m_{N}^{2}m_{a}U_{eN}^{4}}{2\pi f_{a}^{2}}$ (2.25)
@@
No other decay channels are available at tree level. However, as it will be seen below (see section 4), most of the constraints on axion-like particles come from their interactions with electrons and photons. Though the Lagrangian in eq. (2.9) does not produce axion-electron or axion-photon interactions at tree level, such couplings can be induced at 1-loop and 2-loop respectively, as shown in Figure 2.
@@
$g_{a\gamma}$ $\approx\frac{e^{2}g_{ae}}{2\pi^{2}m_{e}}\left(1+\frac{1}{12}\frac{m_{a}^{2}}{m_{e}^{2}}\right)=\frac{\sqrt{2}e^{2}G_{F}|U_{eN}|^{4}m_{N}^{2}}{32\pi^{4}f_{a}}\left(1+\frac{1}{12}\frac{m_{a}^{2}}{m_{e}^{2}}\right),$ (2.26)
@@
Nevertheless, for light ALPs, $m_{a}\lesssim 1$ keV, which is the focus of this study, the only open decay channels for the ALPs are either a pair of active neutrinos or a pair of photons. But, as seen in Fig. 2, the diphoton decay happens at 2-loop and thus it is significantly suppressed with respect to the decay to active neutrinos. For example, the decay width to photons of a 1 keV ALP with decay constant of 1 TeV will be of order $10^{-46}$ GeV, negligible compared to decay width to active neutrinos in eq. (2.25).
@@
$\tau_{a}=1~{}\rm{sec}\times\left(\frac{1~{}\rm{GeV}}{m_{N}}\right)^{2}\times\left(\frac{1~{}\rm{keV}}{m_{a}}\right)\times\left(\frac{2.03\times 10^{-6}}{|U_{eN}|^{2}}\right)^{2}\times\left(\frac{f_{a}}{1~{}\rm{TeV}}\right)^{2}.$ (2.27)
@@ 3 Cosmological history of HNLs and ALPs
The change on the energy and number densities of the various particle species due to the expansion of the universe and the processes in Fig. 3 can be computed with the Boltzmann equations [37, 38]
@@
$\frac{dn_{X}}{dt}+3Hn_{X}$ $=\frac{\delta n_{X}}{\delta t}=\int g_{X}\frac{d^{3}p}{(2\pi)^{3}}\mathcal{C}[f],$ (3.1)
@@
$\gamma_{X\rightarrow Y}=\langle\sigma_{X\rightarrow Y}v\rangle n_{X}^{eq,2}=\langle\Gamma_{X\rightarrow Y}\rangle n_{X}^{eq},$ (3.2)
@@
$\gamma_{X\to Y_{1}\cdots}=n_{X}^{eq}(z)\frac{K_{1}(z)}{K_{2}(z)}\Gamma_{X},$ (3.3)
@@ 3.1 Abundances before BBN
We start by assuming that after some inflationary epoch and subsequent reheating, the SM particles are in equilibrium in the early Universe. The actual value of the reheating temperature is no importance, since most particles will be at their equilibrium densities, but it must be low enough to ensure that the ALPs are not strongly coupled after reheating. The HNLs are thus assumed to be in thermal and kinematical equilibrium with the SM, but the ALPs are not. The evolution of the densities of HNL and ALP can be described by coupled Boltzmann equations, as in eq. (3.1). Before neutrino decoupling, the HNL number density is given by
$\frac{dn_{N}}{dt}+3Hn_{N}=$ $-\left\langle\sigma_{NN\rightarrow SM}~{}v\right\rangle\left(n_{N}^{2}-n_{N}^{\mathrm{eq},2}\right)-\left\langle\sigma_{NN\to aa}~{}v\right\rangle\left(n_{N}^{2}-n_{N}^{\mathrm{eq},2}\frac{n_{a}^{2}}{n_{a}^{\mathrm{eq},2}}\right)$ (3.5)
@@
$\frac{dn_{a}}{dt}+3Hn_{a}$ $=-\left\langle\sigma_{aa\rightarrow\nu\nu}~{}v\right\rangle\left(n_{a}^{2}-n_{a}^{\mathrm{eq},2}\right)-\left\langle\sigma_{aa\to NN}~{}v\right\rangle\left(n_{a}^{2}-n_{a}^{\mathrm{eq},2}\frac{n_{N}^{2}}{n_{N}^{\mathrm{eq},2}}\right)$ (3.6)
@@
$-\gamma_{N\rightarrow SM}\left(\frac{Y_{N}}{Y_{N}^{\mathrm{eq}}}-1\right)-\gamma_{N\rightarrow a\nu}\left(\frac{Y_{N}}{Y_{N}^{\mathrm{eq}}}-\frac{Y_{a}}{Y_{a}^{\mathrm{eq}}}\right),$ (3.8)
@@
$-\gamma_{a\rightarrow\nu\nu}\left(\frac{Y_{a}}{Y_{a}^{\mathrm{eq}}}-1\right)+\gamma_{N\rightarrow a\nu}\left(\frac{Y_{N}}{Y_{N}^{\mathrm{eq}}}-\frac{Y_{a}}{Y_{a}^{\mathrm{eq}}}\right),$ (3.9)
@@ 3.2 Abundance of ALPs after BBN and temperature evolution
As we saw before, in all scenarios we consider, the abundance of HNLs depletes very fast before BBN. As a consequence, the main production mechanisms for ALPs, which were HNL-ALP scattering and HNL decays, are really inefficient after BBN and thus the ALP abundance only decreases at late times. Even the processes $aa\leftrightarrow N\nu$ and $aa\leftrightarrow\nu\nu$ fall out of thermal equilibrium significantly before BBN, as seen below in Fig. 4. Therefore, the abundance of ALPs after the end of BBN is solely determined by their decays. Their Boltzmann equation is therefore given by[12]
$zHs\frac{dY_{a}}{dz}=-\gamma_{a\rightarrow\nu\nu}\left(\frac{Y_{a}}{Y_{a}^{\mathrm{eq}}}-1\right),\quad~{}\text{with}\quad~{}z>z_{\rm BBN}.$ (3.10)
@@
Before neutrino decoupling, the temperatures of the photon and neutrino baths evolved together, changing only via adiabatic cooling since ALP decays are negligible before BBN [39]. After the end of BBN, however, the decays of ALPs to neutrinos, $a\to\nu\nu$ become relevant, which will introduce energy into the neutrino bath and thus increase its temperature. An increase on the neutrino temperature relative with the photon temperature at the time of recombination has a strong impact on the number of effective neutrino degrees of freedom $N_{\rm eff}$. Because of this, we need to track separately the evolution of the photon and neutrino temperatures from the end of BBN until the formation of the CMB. Neglecting chemical potentials and effects from neutrino oscillations, the evolution of the temperatures with time is given by [38, 40]
@@
$\frac{dT_{\nu}}{dt}$ $=-\frac{12H\rho_{\nu}+3H(\rho_{a}+p_{a})+\delta\rho_{a}/\delta t}{3\tfrac{\partial\rho_{\nu}}{\partial T_{\nu}}+\tfrac{\partial\rho_{a}}{\partial T_{\nu}}}$ (3.11)
where $\rho_{i}$ and $p_{i}$ are the energy density and pressure of particle species $i$, and $\delta\rho_{a}/\delta t$ is the energy exchange rate from ALPs to neutrinos, given by the collision operator in eq.(3.1). Equation (3.11) takes into account the decoupling of the electrons from the photon bath when they become non-relativistic, as well as the decoupling of the ALP from the neutrino bath when $T_{\nu}\gtrsim m_{a}$.
@@ 4 Constraints from cosmology, astrophysics and direct searches
In this work we have introduced two new particles, a HNL and an ALP that modify very significantly the evolution of the early Universe compared to $\Lambda$CDM. The cosmological and astrophysical implications of both of these species are well known separately [41, 27, 42]. However, their combination has a somewhat different effect, as the HNLs are allowed to be lighter than in most previous scenarios, due to the weakening of the BBN bound (see below), and also due to the ALPs coupling exclusively to neutrinos via the HNL portal. Consequently we need to evaluate the predictions of our model towards known cosmological and astrophysical observables and hence assess whether their strong constraints render the model invalid or there are still parameter combinations that are unconstrained. For this purpose we study how this model affects the formation of primordial elements (BBN), the observations of the CMB, as well as astrophysical constraints such as those from the observation of the supernova SN1987A, and others such as extra-galactic background light (EBL) or X-ray constraints. Additionally, the new decay channel for HNLs available in this model modifies the prospects of direct searches for HNLs at colliders, fixed target or beam dump experiments. Consequently, we also study the sensitivity of our model to direct searches for HNLs at future facilities.
@@ 4.1 Big Bang Nucleosynthesis
The predictions of $\Lambda$CDM with regards to the formation of the primordial elements match very well with their observed present abundances [^2] The present-day abundance of ${}^{7}Li$ is currently in disagreement with the predictions from standard cosmology. Attempts have been made to provide an explanation in modified cosmologies, with varying degrees of success [43, 44]. Therefore, it remains unclear whether this is caused by modifications over $\Lambda$CDM or an inaccurate measurement of their present abundance. [45, 46]. These predictions are, however very sensitive to the cosmological state of the Universe at any time between neutrino decoupling and the CMB. Any modification of the temperature of neutrino decoupling, the rate of expansion of the Universe or energy injection in the primordial plasma, may cause a catastrophic change on the formation of light elements. This is especially true in our scenario, where the decays of HNLs into mesons disturb the $p\leftrightarrow n$ conversion processes that set the initial proton and neutron abundances for BBN [47], other works in this direction can be found in [48, 49, 50, 51, 52]. If the HNLs decay fast enough, however, the abundances of protons and neutrons have time to restore to the expected values from $\Lambda$CDM, and thus there is no effect on BBN.
Figure 5: Upper limit on the HNL lifetime $\tau_{N}$ as a function of its mass $m_{N}$ by requiring that the HNL decay products do not alter BBN significantly. The plot is adapted from [47], for the case with no ALP (dashed blue) and two scenarios with $f_{a}=1$ TeV (magenta) and $f_{a}=10^{2.5}$ TeV (green), respectively.
Such considerations can be used to set an upper limit on the HNL lifetime such that it decays sufficiently early without disturbing the BBN abundances. In Fig. 5, we show the upper limit on the HNL lifetime as a function of its mass. The blue dashed curve corresponds to the typically considered scenario with an HNL only, while the magenta and green solid curves represent scenarios with the ALP and the decay constant $f_{a}=1$ TeV and $f_{a}=10^{2.5}$ TeV, respectively. The limit is adapted from [47] and it is comparable to earlier results in [53, 54, 55]. In the HNL + ALP scenarios, the decay of HNLs to mesons is suppressed, which relaxes the constraint for $m_{N}\gtrsim m_{\pi}$.
@@
In addition to the decays of the HNLs, the decays of ALPs may also modify the primordial plasma and affect BBN. If ALPs decay before neutrino decoupling, since their primary decay channel is $a\to\nu\nu$, they would modify the neutrino spectrum and increase the temperature of the neutrino bath, thus delaying neutrino decoupling and the formation of primordial elements. Consequently, we do not consider scenarios where the ALPs have short lifetimes and to ensure that they do not directly affect BBN, we only consider ALPs that decay after the end of BBN, that is, $\tau_{a}>10^{4}$  s. Subsequent decays of ALPs may also dissociate the formed elements and modify their abundance [39], however, this dissociation is mostly caused by electromagnetic cascades, which are rare in our scenario, since the decay rate of $a\to\gamma\gamma$ is negligible. In general, late decays of long-lived particles may in principle affect the abundance of light elements through high-energy electromagnetic and hadronic cascades [56, 49, 57, 58, 59, 60]. Even though our ALP only decays dominantly to neutrinos, cascades and scattering with background neutrinos can dissociate elements, but corresponding constraints apply to heavy long-lived particles whereas we consider light ALPs, $m_{a}\lesssim 1$ keV that are non-relativistic at the time of decay. The produced neutrinos with kinetic energies $\lesssim 1$ keV are not expected to have any significant impact.
@@
On the other hand, since ALPs are relativistic before neutrino decoupling, a sufficient abundance would act as dark radiation and thus modify the Hubble rate during radiation domination. This effect can be understood as increase in the neutrino temperature and therefore results in a modification on the value of $N_{\rm eff}$ at the time of BBN [13]. Since the contribution from dark radiation is only relevant when the ALPs are close to thermal equilibrium, we can parametrize the deviation from $\Lambda$CDM as
$\Delta N_{\rm eff}^{\rm BBN}\approx\frac{\rho_{a}}{\rho_{\gamma}}\approx\frac{\rho_{a}^{eq}n_{a}}{\rho_{\gamma}n_{a}^{eq}}$ (4.1)
@@
where $\rho_{\gamma}$ is the energy density in photons at BBN. Though $N_{\rm eff}$ is not directly measured at BBN, it can inferred from the measured value at recombination by Planck [61], to be [62] $N_{\rm eff}^{\rm BBN}=2.86\pm 0.15$, hence we set an upper bound on the ALP abundance at BBN by requiring $\Delta N_{\rm eff}^{\rm BBN}\lesssim 0.2$.